OpenStudy (anonymous):

Which is the equation of the line, in point-slope form, that has a slope of –4 and passes through the point (7, 1)? y – 7 = 4(x – 1) y – 1 = –4(x + 7) y – 1 = –4(x – 7) y – 7 = –4(x – 1)

OpenStudy (anonymous):

C

TheSmartOne (thesmartone):

\[(y-y _{1}=m(x-x _{1})\]

TheSmartOne (thesmartone):

Here \[(x _{1},y _{1}) is (7,1)\]

TheSmartOne (thesmartone):

and m=-4

TheSmartOne (thesmartone):

m is nothing but the slope

TheSmartOne (thesmartone):

@twood03

TheSmartOne (thesmartone):

So your answer is C.

TheSmartOne (thesmartone):

(y-7)=-4(x-7)

TheSmartOne (thesmartone):

When you substitute everything you will get (y-7)=-4(x-7)